OpenStudy (anonymous):

use the distance formula to find th distance between each pair of points. a. (3,7), (7,7) b. (9,2), (10,4) c.(5,3), (8,3)

OpenStudy (anonymous):

For a pair of points, \[(x_1,y_1),(x_2,y_2)\]the distance is\[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]For the first pair:\[d_1=\sqrt{(7-3)^2+(7-7)^2}=\sqrt{4^2}=4\]For the second pair:\[d_2=\sqrt{(10-9)^2+(4-2)^2}=\sqrt{1+4}=\sqrt{5}\]For the third:\[d_3=\sqrt{(8-5)^2+(3-3)^2}=\sqrt{3^2}=3\]

OpenStudy (anonymous):

great thanks Lokisan :)

OpenStudy (anonymous):

np